### v3

# Limit Filters and Dependent Choice in Countable-Support Symmetric Iterations

Abstract
We isolate the limit-stage filter construction needed for countable-support symmetric iterations built from standard successor-step symmetric systems. At successor stages we take the $\omega_{1}$-completion of the usual successor-stage symmetry filter. At limit stages of uncountable cofinality, countable supports are bounded and the direct-limit filter is $\omega_{1}$-complete by stage-bounding; at limits of cofinality $\omega$, we define the limit filter as the smallest normal $\omega_{1}$-complete filter extending the head-pullback generators. In all cases the resulting limit filter is normal and $\omega_{1}$-complete.
Using these limit filters, we prove that the class of hereditarily symmetric names is closed under the operations required for $\mathrm{ZF}$, that the resulting symmetric model satisfies $\mathrm{ZF}$ (Theorem 3.17), and that over a $\mathrm{ZFC}$ ground it satisfies $\mathrm{DC}=\mathrm{DC}_{\omega}$ (Theorem 3.20). We also explain why no general class-length $\mathrm{ZF}$-preservation theorem is claimed here, although specific class-length symmetric iterations may still be analyzable separately over a $\mathrm{GBC}$ background.
As a self-contained application, we construct, for any uncountable cardinal $\kappa$ with $\mathrm{cf}(\kappa)\geq\omega_{1}$, a model of $\mathrm{ZF}+\mathrm{DC}+\neg\mathrm{AC}_{\kappa}(\mathcal{F})$, where $\mathcal{F}$ is a $\kappa$-indexed family of $2$-element sets of reals with no choice function. Each step of the iteration adds one unpickable pair, so the degree of $\mathrm{AC}$-failure is controlled by the iteration length. We also prove that the analogous finite-support construction fails $\mathrm{DC}$ at the first $\omega$-limit stage, showing why $\omega_{1}$-complete limit filters are structurally needed for this application.
1 Introduction
Symmetric extensions produce intermediate models of $\mathrm{ZF}$ between a ground model $V$ and a forcing extension $V[G]$ by restricting attention to the class of hereditarily symmetric names. Karagila developed a general framework for finite-support symmetric iterations [3]. Countable support introduces a new constraint at limit stages of cofinality $\omega$: to keep the class of hereditarily symmetric names closed under countable constructions, one needs limit-stage symmetry filters which are not merely normal but also $\omega_{1}$–complete.
What is proved here.
Working in a fixed background universe $V$, we assume the standard successor-step construction of symmetric systems (as in [3, §§3–5]) and isolate the additional limit-stage technology needed for countable-support iterations. At each limit $\lambda$ we define a canonical limit filter $\tilde{\mathcal{F}}_{\lambda}$ extending the head pullbacks of earlier-stage filters and prove:
  • $\tilde{\mathcal{F}}_{\lambda}$ is normal and $\omega_{1}$–complete (Theorem 3.13);
  • the class $\mathrm{HS}_{\lambda}$ of hereditarily symmetric names is closed under the operations required in the standard symmetric-model verification of $\mathrm{ZF}$ (Theorems 3.15 and 3.17);
  • if $V\models\mathrm{ZFC}$, then the resulting symmetric model satisfies $\mathrm{DC}=\mathrm{DC}_{\omega}$ (Theorem 3.20).
Why the limit filters matter.
The central point is that countable-support iterations require $\omega_{1}$-complete symmetry filters at limit stages. This has two consequences needed in the present paper. First, $\omega_{1}$-completeness ensures that hereditarily symmetric names are closed under countable tuples (Theorem 3.15), yielding $\mathrm{DC}$ in the resulting model (Theorem 3.20) and ruling out Dedekind-finite or amorphous pathologies (Corollary 3.22). Second, the same limit-filter structure is what makes the later application possible: the iterated unordered-pairs construction of Section 4 uses these filters to build, for each uncountable $\kappa$ with $\mathrm{cf}(\kappa)\geq\omega_{1}$, a model of $\mathrm{ZF}+\mathrm{DC}+\neg\mathrm{AC}_{\kappa}(\mathcal{F})$.
A finite-support version of the same step forcing fails to produce a $\mathrm{DC}$ model at the $\omega$-th limit stage (Proposition 4.4), showing that the $\omega_{1}$-complete limit-filter construction is not merely convenient but structurally necessary for this application.
What is not claimed.
We do not claim here a complete general theory of countable-support symmetric iterations comparable in scope to broader support frameworks. The successor-step construction is taken from the standard theory, and the new contribution of the present paper is the limit-stage filter technology needed for countable support together with the resulting $\mathrm{ZF}/\mathrm{DC}$ preservation theorems.
We also do not claim that the framework by itself produces new choiceless consequences; those arise from the specific iteration templates used in applications (cf. Section 4). In particular, we do not claim any uncountable-length dependent choice (e.g. $\mathrm{DC}_{\omega_{1}}$) without additional hypotheses, and we do not address cardinal preservation or other fine structure beyond what is explicitly proved. All uses of “countable” refer to countability in the background universe $V$ (Remark 2.1); for the only countability closure needed here (unions of countable supports, i.e. subsets of ordinals), see Remark 2.2.
Organization.
Section 2 fixes forcing/symmetry notation and records the Symmetry Lemma and related invariance facts used later. Section 3 develops the countable-support iteration, the limit filters, and the closure and preservation theorems. Remark 3.25 isolates the points where finite support is insufficient for the present arguments. Section 4 constructs the iterated unordered pairs model as a self-contained application, producing $\mathrm{ZF}+\mathrm{DC}+\neg\mathrm{AC}_{\kappa}(\mathcal{F})$ for any uncountable $\kappa$ with $\mathrm{cf}(\kappa)\geq\omega_{1}$, and proves that the finite-support version of the same step forcing fails to preserve $\mathrm{DC}$.
2 Preliminaries: forcing, automorphisms, and symmetric names
2.1 Background theory and metatheoretic conventions
Remark 2.1 (Metatheory vs. object theory).
All constructions of forcing notions, names, automorphism groups, filters, and supports are carried out in a fixed background universe $V$.
  • For the set-length development (existence of the iteration, the limit filters, and $\mathrm{ZF}$ preservation) we work in $V\models\mathsf{ZF}$.
  • For the dependent-choice conclusion (Theorem 3.20) we additionally assume $V\models\mathsf{ZFC}$.
When we say a set is countable, this is always in the sense of the background universe $V$ (equivalently: there is an injection into $\omega$ in $V$).
No claim is made that these sets are countable inside intermediate symmetric models.
Remark 2.2 (Countable unions of countable supports in $\mathrm{ZF}$).
All countability bookkeeping in Sections 33.5 concerns subsets of ordinals (supports and stage indices). In $\mathrm{ZF}$, if $\langle A_{n}:n<\omega\rangle$ is a sequence of countable subsets of an ordinal $\lambda$, then $\bigcup_{n<\omega}A_{n}$ is countable. Indeed, each $A_{n}\subseteq\lambda$ admits a canonical increasing enumeration (defined by recursion using the well-ordering of $\lambda$, padded by 0 if $A_{n}$ is finite), and then $(n,m)\mapsto e_{n}(m)$ is a surjection from $\omega\times\omega$ onto $\bigcup_{n<\omega}A_{n}$. Thus no separate choice principle about countable unions of arbitrary sets is used for the set-length iteration technology here.
2.2 Forcing and names
We use standard forcing conventions. A forcing notion is a poset $\mathbb{P}=(P,\leq)$ with top condition $1_{\mathbb{P}}$. A $\mathbb{P}$-name is a set $\dot{x}$ of pairs $\langle\dot{y},p\rangle$ with $\dot{y}$ a $\mathbb{P}$-name and $p\in\mathbb{P}$. For a $V$-generic filter $G\subseteq\mathbb{P}$, $\dot{x}^{G}$ denotes the usual valuation.
2.3 Automorphisms and symmetric systems
Definition 2.3 (Automorphisms).
An automorphism of $\mathbb{P}$ is a bijection $\pi:P\to P$ satisfying $p\leq q\iff\pi(p)\leq\pi(q)$ and $\pi(1_{\mathbb{P}})=1_{\mathbb{P}}$. If $\mathcal{G}\leq\mathrm{Aut}(\mathbb{P})$ is a subgroup, we write $\pi\cdot p$ for $\pi(p)$.
Definition 2.4 (Normal filter of subgroups).
Let $\mathcal{G}$ be a group. A filter of subgroups $\mathcal{F}$ on $\mathcal{G}$ is a nonempty collection of subgroups of $\mathcal{G}$ such that:
  1. 1.
    $\mathcal{G}\in\mathcal{F}$;
  2. 2.
    if $H\in\mathcal{F}$ and $H\leq K\leq\mathcal{G}$, then $K\in\mathcal{F}$;
  3. 3.
    if $H,K\in\mathcal{F}$, then $H\cap K\in\mathcal{F}$.
It is normal if for all $H\in\mathcal{F}$ and all $\pi\in\mathcal{G}$, $\pi H\pi^{-1}\in\mathcal{F}$. It is $\omega_{1}$-complete if $\bigcap_{n<\omega}H_{n}\in\mathcal{F}$ whenever $H_{n}\in\mathcal{F}$ for all $n<\omega$.
Definition 2.5 ($\omega_{1}$-completion of a normal filter).
Let $\mathcal{F}$ be a normal filter of subgroups on $\mathcal{G}$. Define its $\omega_{1}$-completion $\mathcal{F}^{\omega_{1}}$ by
$K\in\mathcal{F}^{\omega_{1}}\quad\Longleftrightarrow\quad\exists\langle H_{n}:n<\omega\rangle\in\mathcal{F}^{\omega}\ \Bigl(\ \bigcap_{n<\omega}H_{n}\leq K\ \Bigr).$
Lemma 2.6 ($\omega_{1}$-completion is the smallest normal $\omega_{1}$-complete extension).
If $\mathcal{F}$ is a normal filter of subgroups on $\mathcal{G}$, then $\mathcal{F}^{\omega_{1}}$ is a normal, $\omega_{1}$-complete filter of subgroups on $\mathcal{G}$, and it is the smallest normal, $\omega_{1}$-complete filter extending $\mathcal{F}$.
Proof.
Upward closure and containing $\mathcal{G}$ are immediate. $\omega_{1}$-completeness holds because if $K_{m}\in\mathcal{F}^{\omega_{1}}$ is witnessed by $\langle H^{m}_{n}:n<\omega\rangle\subseteq\mathcal{F}$, then $\bigcap_{m}K_{m}$ is witnessed by the single $\mathcal{F}$-sequence obtained by concatenating the sequences $\langle H^{m}_{n}:n<\omega\rangle$ over $m<\omega$. Normality is preserved since for $\pi\in\mathcal{G}$,
$\pi\Bigl(\bigcap_{n<\omega}H_{n}\Bigr)\pi^{-1}=\bigcap_{n<\omega}\pi H_{n}\pi^{-1},$
and $\mathcal{F}$ is normal. Minimality is immediate: any normal $\omega_{1}$-complete filter extending $\mathcal{F}$ must contain every countable intersection of members of $\mathcal{F}$, hence must contain $\mathcal{F}^{\omega_{1}}$. ∎
Definition 2.7 (Symmetric system).
A symmetric system is a triple $\langle\mathbb{P},\mathcal{G},\mathcal{F}\rangle$ where $\mathbb{P}$ is a forcing notion, $\mathcal{G}\leq\mathrm{Aut}(\mathbb{P})$, and $\mathcal{F}$ is a normal filter of subgroups of $\mathcal{G}$.
2.4 Action on names and the Symmetry Lemma
Given $\langle\mathbb{P},\mathcal{G},\mathcal{F}\rangle$, define recursively the action of $\pi\in\mathcal{G}$ on $\mathbb{P}$-names by
$\pi\dot{x}\ :=\ \{\langle\pi\dot{y},\ \pi p\rangle:\langle\dot{y},p\rangle\in\dot{x}\}.$
Definition 2.8 (Stabilizers).
For a name $\dot{x}$, its (setwise) stabilizer is
$\operatorname{sym}_{\mathcal{G}}(\dot{x})\ :=\ \{\pi\in\mathcal{G}:\pi\dot{x}=\dot{x}\}.$
When $\mathcal{G}$ is clear from context, we write $\operatorname{sym}(\dot{x})$.
Lemma 2.9 (Symmetry Lemma / Permutation Lemma).
For any formula $\varphi$ in the forcing language, any $\pi\in\mathcal{G}$, any $p\in\mathbb{P}$, and any names $\dot{x}_{1},\dots,\dot{x}_{n}$,
$p\Vdash_{\mathbb{P}}\varphi(\dot{x}_{1},\dots,\dot{x}_{n})\quad\Longleftrightarrow\quad\pi p\Vdash_{\mathbb{P}}\varphi(\pi\dot{x}_{1},\dots,\pi\dot{x}_{n}).$
See, e.g., [3, Lemma 2.4].
2.5 Symmetric and hereditarily symmetric names
Definition 2.10 (Symmetric / hereditarily symmetric).
Let $\langle\mathbb{P},\mathcal{G},\mathcal{F}\rangle$ be a symmetric system. A name $\dot{x}$ is $\mathcal{F}$-symmetric if $\operatorname{sym}_{\mathcal{G}}(\dot{x})\in\mathcal{F}$. A name $\dot{x}$ is hereditarily $\mathcal{F}$-symmetric if $\dot{x}$ is $\mathcal{F}$-symmetric and every $\dot{y}$ appearing in the transitive closure of $\dot{x}$ (equivalently, every $\dot{y}\in\mathrm{dom}(\dot{x})$ and recursively) is hereditarily $\mathcal{F}$-symmetric. We write $\mathrm{HS}(\mathbb{P},\mathcal{G},\mathcal{F})$ for the class of hereditarily $\mathcal{F}$-symmetric $\mathbb{P}$-names.
2.6 Dependent Choice (set formulation)
Definition 2.11 (Dependent Choice).
Let $\mu$ be a nonzero ordinal. $\mathrm{DC}_{\mu}$ is the statement: for every set $A$ and every binary relation $R\subseteq A\times A$ such that $\forall x\in A\,\exists y\in A\ (xRy)$, there exists a sequence $\langle a_{\xi}:\xi<\mu\rangle$ of elements of $A$ with $a_{\xi}Ra_{\xi+1}$ for all $\xi+1<\mu$. We write $\mathrm{DC}=\mathrm{DC}_{\omega}$ and $\mathrm{DC}_{<\kappa}$ for $\forall\mu<\kappa\,\mathrm{DC}_{\mu}$.
2.7 Two-step iterations (factorization)
Fact 2.12 (Two-step factorization).
[…]
Remark 2.13 (Notation discipline for later sections).
In the iteration sections we will not overload projection symbols. For forcing, we use restriction $p\upharpoonright\beta$ and (when needed) a map $\pi_{\beta,\lambda}:\mathbb{P}_{\lambda}\to\mathbb{P}_{\beta}$. For groups, we use a distinct restriction homomorphism $\rho_{\beta,\lambda}:\mathcal{G}_{\lambda}\to\mathcal{G}_{\beta}$.
3 Limit-Stage Filters for Countable-Support Symmetric Iterations
3.1 Framework and Setup
Definition 3.1 (Countable-support iteration setup).
Fix an ordinal $\Theta$. A countable-support symmetric iteration of length $\Theta$ is a system
$\langle(\mathbb{P}_{\alpha},\mathcal{G}_{\alpha},\mathcal{F}_{\alpha}):\alpha\leq\Theta\rangle$
together with, for each $\alpha<\Theta$, a $\mathbb{P}_{\alpha}$-name
$\dot{\mathscr{S}}_{\alpha}=\langle\dot{\mathbb{Q}}_{\alpha},\dot{\mathcal{H}}_{\alpha},\dot{\mathcal{K}}_{\alpha}\rangle$
such that
$1_{\mathbb{P}_{\alpha}}\Vdash_{\mathbb{P}_{\alpha}}\text{``$\dot{\mathscr{S}}_{\alpha}$ is a symmetric system.''}$
The recursion is as follows.
Stage 0.
$\mathbb{P}_{0}=\{1\}$, $\mathcal{G}_{0}=\{\mathrm{id}\}$, and $\mathcal{F}_{0}=\{\mathcal{G}_{0}\}$.
Successor stage $\alpha+1$.
Let $\dot{\mathscr{S}}_{\alpha}=\langle\dot{\mathbb{Q}}_{\alpha},\dot{\mathcal{H}}_{\alpha},\dot{\mathcal{K}}_{\alpha}\rangle$ be as above. Define
$\mathbb{P}_{\alpha+1}=\mathbb{P}_{\alpha}*\dot{\mathbb{Q}}_{\alpha}.$
Define $\mathcal{G}_{\alpha+1}$ and an intermediate normal successor-stage filter $\mathcal{F}^{\mathrm{Kar}}_{\alpha+1}$ by the standard two-step symmetric-system construction from $(\mathbb{P}_{\alpha},\mathcal{G}_{\alpha},\mathcal{F}_{\alpha})$ and the name $\dot{\mathscr{S}}_{\alpha}$: there is a canonical restriction homomorphism (“head restriction”)
$\rho_{\alpha,\alpha+1}:\mathcal{G}_{\alpha+1}\to\mathcal{G}_{\alpha}$
and $\mathcal{F}^{\mathrm{Kar}}_{\alpha+1}$ is a normal filter of subgroups of $\mathcal{G}_{\alpha+1}$ extending the pullbacks $\rho_{\alpha,\alpha+1}^{-1}(H)$ for $H\in\mathcal{F}_{\alpha}$ and the tail symmetries coming from $\dot{\mathcal{K}}_{\alpha}$. (We do not reprove this successor-step construction here; see, e.g., [3, §§3–5].)
For the countable-support limit-stage arguments below, we require $\omega_{1}$–completeness at intermediate stages, so we define
$\mathcal{F}_{\alpha+1}\ :=\ \bigl(\mathcal{F}^{\mathrm{Kar}}_{\alpha+1}\bigr)^{\omega_{1}},$
the $\omega_{1}$–completion of $\mathcal{F}^{\mathrm{Kar}}_{\alpha+1}$ in the sense of Definition 2.5 (which is the smallest normal $\omega_{1}$–complete extension by Lemma 2.6).
Limit stage $\lambda\leq\Theta$ with $\mathrm{cf}(\lambda)=\omega$.
$\mathbb{P}_{\lambda}$ is the usual countable-support limit of $\langle\mathbb{P}_{\alpha},\dot{\mathbb{Q}}_{\alpha}:\alpha<\lambda\rangle$: conditions are partial functions $p$ with countable domain $\operatorname{dom}(p)\subseteq\lambda$, where for each $\beta\in\operatorname{dom}(p)$, $p(\beta)$ is a $\mathbb{P}_{\beta}$-name for a condition in $\dot{\mathbb{Q}}_{\beta}$, ordered by $q\leq p$ iff $\operatorname{dom}(q)\supseteq\operatorname{dom}(p)$ and for all $\beta\in\operatorname{dom}(p)$,
$q\upharpoonright\beta\ \Vdash_{\mathbb{P}_{\beta}}\ q(\beta)\leq p(\beta).$
Define $\mathrm{supp}(p):=\operatorname{dom}(p)$.
$\mathcal{G}_{\lambda}$ is the inverse limit of $\langle\mathcal{G}_{\beta},\rho_{\gamma,\beta}:\gamma<\beta<\lambda\rangle$ and acts coordinatewise on $\mathbb{P}_{\lambda}$ (Lemma 3.3). The limit filter $\tilde{\mathcal{F}}_{\lambda}$ is defined in Definition 3.9.
Limit stage $\lambda\leq\Theta$ with $\mathrm{cf}(\lambda)\geq\omega_{1}$.
We identify the countable-support limit with the direct limit (Remark 3.2):
$\mathbb{P}_{\lambda}=\bigcup_{\beta<\lambda}\mathbb{P}_{\beta},\qquad\mathcal{G}_{\lambda}=\bigcup_{\beta<\lambda}\mathcal{G}_{\beta},$
with the canonical restriction maps $\pi_{\beta,\lambda}(p)=p\upharpoonright\beta$ and $\rho_{\beta,\lambda}$. The limit filter $\tilde{\mathcal{F}}_{\lambda}$ is the normal filter generated by head pullbacks from earlier stages (Definition 3.9), and it is $\omega_{1}$-complete by stage-bounding (Remark 3.5).
Remark 3.2 (Bookkeeping at uncountable cofinality).
Let $\lambda$ be a limit ordinal with $\mathrm{cf}(\lambda)\geq\omega_{1}$, and let $\mathbb{P}_{\lambda}$ be the usual countable-support limit of $\langle\mathbb{P}_{\alpha},\dot{\mathbb{Q}}_{\alpha}:\alpha<\lambda\rangle$. Then every $p\in\mathbb{P}_{\lambda}$ has bounded support: since $\mathrm{supp}(p)$ is countable and $\mathrm{cf}(\lambda)>\omega$, there is $\beta<\lambda$ with $\mathrm{supp}(p)\subseteq\beta$, hence $p\in\mathbb{P}_{\beta}$. Therefore $\mathbb{P}_{\lambda}=\bigcup_{\beta<\lambda}\mathbb{P}_{\beta}$ (direct limit identification), and similarly $\mathcal{G}_{\lambda}=\bigcup_{\beta<\lambda}\mathcal{G}_{\beta}$ with the evident restriction maps.
Lemma 3.3 (Limit-stage coordinatewise action).
Let $\lambda$ be a limit stage with $\mathrm{cf}(\lambda)=\omega$. If $g=\langle g_{\beta}:\beta<\lambda\rangle\in\mathcal{G}_{\lambda}$, define $(g\cdot p)(\beta)=g_{\beta}\cdot p(\beta)$ for $\beta\in\operatorname{dom}(p)$. Then $p\mapsto g\cdot p$ is an automorphism of $\mathbb{P}_{\lambda}$.
Proof.
It is immediate that $\operatorname{dom}(g\cdot p)=\operatorname{dom}(p)$ and $g\cdot 1_{\mathbb{P}_{\lambda}}=1_{\mathbb{P}_{\lambda}}$. For order preservation, suppose $q\leq p$ in $\mathbb{P}_{\lambda}$. Fix $\beta\in\operatorname{dom}(p)$. Then $q\upharpoonright\beta\Vdash_{\mathbb{P}_{\beta}}q(\beta)\leq p(\beta)$. Apply the invariance of the forcing relation under automorphisms from Section 2 (Lemma 2.9 in its forcing-theorem form): since $g\upharpoonright\beta\in\mathcal{G}_{\beta}$,
$(g\cdot q)\upharpoonright\beta\ \Vdash_{\mathbb{P}_{\beta}}\ (g\cdot q)(\beta)\leq(g\cdot p)(\beta).$
As this holds for all $\beta\in\operatorname{dom}(p)$, we have $g\cdot q\leq g\cdot p$. Bijectivity follows from applying the same argument to $g^{-1}$. ∎
Remark 3.4 (Comparison with related frameworks).
The successor-step symmetric-system construction used above is standard (see, e.g., [3, §§3–5]). What is new in the present paper is the limit-stage treatment needed for countable support: at $\mathrm{cf}(\lambda)\geq\omega_{1}$ the direct-limit identification (Remark 3.2) allows stage-bounding, while at $\mathrm{cf}(\lambda)=\omega$ we define the limit filter by taking the smallest normal $\omega_{1}$-complete extension of the head pullbacks (Definition 3.9).
The independent concurrent work of Karagila–Schilhan [6] develops a broad structural theory of symmetric extensions—including products, reduced products, quotients, and a general $I$-support iteration framework (where $I$ is any ideal on the index ordinal)—and proves results on Kinna–Wagner principles and the HOD theorem. Their $I$-support definition subsumes countable support as a special case. The present paper is complementary in focus: we prove that the specific $\omega_{1}$-complete normal limit filter construction at cofinality-$\omega$ stages (Theorem 3.13) yields DC preservation (Theorem 3.20), which is not addressed in [6]. Accordingly, the present paper should be read as a limit-stage supplement to the standard successor-step theory, rather than as a replacement for more general support frameworks.
Remark 3.5 (Uncountable cofinality: stage-bounding yields $\omega_{1}$-completeness).
Assume $\mathrm{cf}(\lambda)\geq\omega_{1}$ and work in the direct-limit identification of Remark 3.2. Let $\tilde{\mathcal{F}}_{\lambda}$ be the limit filter from Definition 3.9(1), i.e. the smallest normal filter on $\mathcal{G}_{\lambda}$ containing the head pullbacks $\rho_{\beta,\lambda}^{-1}(H)$ for $\beta<\lambda$ and $H\in\mathcal{F}_{\beta}$.
Then $\tilde{\mathcal{F}}_{\lambda}$ is $\omega_{1}$-complete. Indeed, let $\langle K_{n}:n<\omega\rangle\subseteq\tilde{\mathcal{F}}_{\lambda}$. For each $n$, by the normal-generation clause in Definition 3.9(1), fix witnesses $k_{n}<\omega$, ordinals $\beta_{n,i}<\lambda$, groups $H_{n,i}\in\mathcal{F}_{\beta_{n,i}}$, and elements $g_{n,i}\in\mathcal{G}_{\lambda}$ such that
$\bigcap_{i<k_{n}}g_{n,i}\,\rho_{\beta_{n,i},\lambda}^{-1}(H_{n,i})\,g_{n,i}^{-1}\ \leq\ K_{n}.$
Since $\mathrm{cf}(\lambda)\geq\omega_{1}$, the set $\{\beta_{n,i}:n<\omega,\ i<k_{n}\}$ is countable and hence bounded in $\lambda$; let $\beta^{*}<\lambda$ bound it.
For each $n,i$, set $H^{\prime}_{n,i}:=\rho_{\beta_{n,i},\beta^{*}}^{-1}(H_{n,i})$. Then $H^{\prime}_{n,i}\in\mathcal{F}_{\beta^{*}}$ by Lemma 3.7, and
$\rho_{\beta^{*},\lambda}^{-1}(H^{\prime}_{n,i})\ \leq\ \rho_{\beta_{n,i},\lambda}^{-1}(H_{n,i})$
by coherence of restrictions. Replacing each head pullback by $\rho_{\beta^{*},\lambda}^{-1}(H^{\prime}_{n,i})$ only strengthens the left-hand side, so we may assume all generators come from stage $\beta^{*}$.
Now use $\omega_{1}$-completeness of $\mathcal{F}_{\beta^{*}}$ (by induction on stages, using the successor-stage $\omega_{1}$-completion) to form the countable intersection of the corresponding (conjugated) stage-$\beta^{*}$ generators inside $\mathcal{F}_{\beta^{*}}$, and pull back along $\rho_{\beta^{*},\lambda}$. This produces an element of $\tilde{\mathcal{F}}_{\lambda}$ contained in $\bigcap_{n<\omega}K_{n}$.
3.2 Filter Generation at Limit Stages
Let $\lambda$ be a limit ordinal and assume inductively that each $\mathcal{F}_{\beta}$ for $\beta<\lambda$ is a normal, $\omega_{1}$-complete filter on $\mathcal{G}_{\beta}$. We define the limit filter $\tilde{\mathcal{F}}_{\lambda}$ uniformly from the head pullbacks $\rho_{\beta,\lambda}^{-1}(H)$ (Definition 3.9). At $\mathrm{cf}(\lambda)\geq\omega_{1}$ we take the smallest normal filter containing these generators, and $\omega_{1}$-completeness follows by stage-bounding (Remark 3.5). At $\mathrm{cf}(\lambda)=\omega$ we instead take the smallest normal $\omega_{1}$-complete filter extending the same generator family.
Remark 3.6 (Head pullbacks of subgroups).
For $\beta<\lambda$ and $H\leq\mathcal{G}_{\beta}$, we write
$\rho_{\beta,\lambda}^{-1}(H)=\{g\in\mathcal{G}_{\lambda}:\rho_{\beta,\lambda}(g)\in H\}.$
When pulling back filters along homomorphisms we use the notation $\pi^{*}\mathcal{F}$ from Lemma 3.7.
Lemma 3.7 (Pullbacks and restrictions preserve normality and $\omega_{1}$-completeness).
Let $\pi:\mathcal{G}\to\mathcal{H}$ be a homomorphism and let $\mathcal{F}$ be a normal, $\omega_{1}$-complete filter on $\mathcal{H}$. Define the pullback filter on $\mathcal{G}$ by
$\pi^{*}\mathcal{F}\ :=\ \{\,K\leq\mathcal{G}:\ \exists H\in\mathcal{F}\ \text{with}\ \pi^{-1}(H)\leq K\,\}.$
Then $\pi^{*}\mathcal{F}$ is a normal, $\omega_{1}$-complete filter on $\mathcal{G}$.
Moreover, if $\iota:\mathcal{H}\hookrightarrow\mathcal{G}$ is an inclusion and $\mathcal{F}$ is a normal, $\omega_{1}$-complete filter on $\mathcal{G}$, then the restriction
$\iota^{*}\mathcal{F}\ :=\ \{\,K\leq\mathcal{H}:\ \iota(K)\in\mathcal{F}\,\}$
is a normal, $\omega_{1}$-complete filter on $\mathcal{H}$.
Proof.
Pullback. Upward closure and containing $\mathcal{G}$ are immediate. For $\omega_{1}$-completeness, suppose $\langle K_{n}:n<\omega\rangle$ with each $K_{n}\in\pi^{*}\mathcal{F}$. Choose $H_{n}\in\mathcal{F}$ with $\pi^{-1}(H_{n})\leq K_{n}$. Then $\bigcap_{n<\omega}H_{n}\in\mathcal{F}$ by $\omega_{1}$-completeness, and
$\pi^{-1}\!\left(\bigcap_{n<\omega}H_{n}\right)=\bigcap_{n<\omega}\pi^{-1}(H_{n})\ \leq\ \bigcap_{n<\omega}K_{n},$
so $\bigcap_{n<\omega}K_{n}\in\pi^{*}\mathcal{F}$. For normality: if $K\in\pi^{*}\mathcal{F}$ witnessed by $H\in\mathcal{F}$ and $g\in\mathcal{G}$, then $\pi(g)H\pi(g)^{-1}\in\mathcal{F}$ and
$\pi^{-1}(\pi(g)H\pi(g)^{-1})=g\,\pi^{-1}(H)\,g^{-1}\ \leq\ gKg^{-1},$
so $gKg^{-1}\in\pi^{*}\mathcal{F}$.
Restriction. Upward closure and containing $\mathcal{H}$ are immediate. For $\omega_{1}$-completeness, if $K_{n}\in\iota^{*}\mathcal{F}$ for $n<\omega$, then $\iota(K_{n})\in\mathcal{F}$ and
$\iota\!\left(\bigcap_{n<\omega}K_{n}\right)=\bigcap_{n<\omega}\iota(K_{n})\in\mathcal{F},$
so $\bigcap_{n<\omega}K_{n}\in\iota^{*}\mathcal{F}$. Normality is inherited since for $h\in\mathcal{H}$ and $K\in\iota^{*}\mathcal{F}$,
$\iota(hKh^{-1})=h\,\iota(K)\,h^{-1}\in\mathcal{F}.\qed$
Definition 3.8 (Supports of conditions and names (countable-support case)).
Let $\lambda$ be a limit stage.
Support of a condition. For $p\in\mathbb{P}_{\lambda}$, let $\mathrm{supp}(p)\subseteq\lambda$ be the set of coordinates at which $p$ is nontrivial (in the usual sense for the iteration/product presentation at $\lambda$). In the countable-support presentation, $\mathrm{supp}(p)$ is required to be countable.
Support of a name. Define the support of a $\mathbb{P}_{\lambda}$–name $\dot{x}$ by recursion on name-rank:
$\mathrm{supp}(\dot{x})\ :=\ \bigcup\Bigl\{\ \mathrm{supp}(p)\ \cup\ \mathrm{supp}(\sigma)\ :\ (\sigma,p)\in\dot{x}\ \Bigr\}.$
We say that $\dot{x}$ has countable support if $\mathrm{supp}(\dot{x})$ is countable.
Definition 3.9 (Limit filter).
Let $\lambda$ be a limit ordinal and assume we have defined $(\mathbb{P}_{\beta},\mathcal{G}_{\beta},\mathcal{F}_{\beta})$ for all $\beta<\lambda$, with coherent restriction homomorphisms $\rho_{\beta,\lambda}:\mathcal{G}_{\lambda}\to\mathcal{G}_{\beta}$.
Define the head-generator family at $\lambda$ by
$\mathcal{B}_{\lambda}\ :=\ \bigl\{\ \rho_{\beta,\lambda}^{-1}(H)\ :\ \beta<\lambda\ \&\ H\in\mathcal{F}_{\beta}\ \bigr\}.$
  1. 1.
    If $\mathrm{cf}(\lambda)\geq\omega_{1}$, let $\tilde{\mathcal{F}}_{\lambda}$ be the normal filter on $\mathcal{G}_{\lambda}$ generated by $\mathcal{B}_{\lambda}$, equivalently:
    $K\in\tilde{\mathcal{F}}_{\lambda}\ \Longleftrightarrow\ \exists k<\omega\ \exists B_{0},\dots,B_{k-1}\in\mathcal{B}_{\lambda}\ \exists g_{0},\dots,g_{k-1}\in\mathcal{G}_{\lambda}\ \Bigl(\bigcap_{i<k}g_{i}B_{i}g_{i}^{-1}\leq K\Bigr).$
  2. 2.
    If $\mathrm{cf}(\lambda)=\omega$, let $\tilde{\mathcal{F}}_{\lambda}$ be the normal $\omega_{1}$–complete filter on $\mathcal{G}_{\lambda}$ generated by $\mathcal{B}_{\lambda}$, equivalently:
    $K\in\tilde{\mathcal{F}}_{\lambda}\ \Longleftrightarrow\ \exists\langle B_{n}:n<\omega\rangle\in\mathcal{B}_{\lambda}^{\omega}\ \exists\langle g_{n}:n<\omega\rangle\in\mathcal{G}_{\lambda}^{\omega}\ \Bigl(\bigcap_{n<\omega}g_{n}B_{n}g_{n}^{-1}\leq K\Bigr).$
Remark 3.10 (Syntactic support vs. background countability).
By Definition 3.8, $\mathrm{supp}(\dot{x})$ is computed recursively from the supports of the conditions appearing in $\dot{x}$ (and from the supports of subnames). When we say $\dot{x}$ has countable support” in the countable-support part of the paper, we mean that $\mathrm{supp}(\dot{x})$ is countable in the ambient background universe (cf. Remark 2.1).
The $\omega_{1}$–completeness argument splits by cofinality. If $\mathrm{cf}(\lambda)\geq\omega_{1}$, stage-bounding reduces countable intersections in $\tilde{\mathcal{F}}_{\lambda}$ to a single stage filter (Remark 3.5). If $\mathrm{cf}(\lambda)=\omega$, $\omega_{1}$–completeness is built into the definition of $\tilde{\mathcal{F}}_{\lambda}$ as the smallest normal $\omega_{1}$–complete filter containing the head generators (Definition 3.9).
Lemma 3.11 (Stage-Bounding Lemma—uncountable cofinality only).
If $\lambda$ is a limit ordinal with $\mathrm{cf}(\lambda)\geq\omega_{1}$ and $\{\beta_{n}:n<\omega\}\subset\lambda$ is any countable set of ordinals below $\lambda$, then
$\sup_{n<\omega}\beta_{n}\;<\;\lambda.$
Proof.
If $\sup_{n}\beta_{n}=\lambda$, then $\{\beta_{n}:n<\omega\}$ would be a countable cofinal subset of $\lambda$, contradicting $\mathrm{cf}(\lambda)\geq\omega_{1}$. Hence $\sup_{n}\beta_{n}<\lambda$. ∎
Remark 3.12 (Failure of stage-bounding at countable cofinality).
When $\mathrm{cf}(\lambda)=\omega$, the stage-bounding lemma fails. For example, if $\lambda=\omega$, the set $\{n:n<\omega\}$ is countable but $\sup_{n}n=\omega=\lambda$. Similarly, if $\lambda=\aleph_{\omega}$, the sequence $\{\aleph_{n}:n<\omega\}$ is countable and cofinal. This is precisely why, at limits of countable cofinality, $\omega_{1}$–completeness must be built into the limit-filter definition (Definition 3.9(2)), rather than proved by stage-bounding.
Theorem 3.13 (Limit filter properties).
Let $\lambda$ be a limit stage and let $\tilde{\mathcal{F}}_{\lambda}$ be as in Definition 3.9. Assume inductively that each $\mathcal{F}_{\beta}$ for $\beta<\lambda$ is a normal $\omega_{1}$–complete filter on $\mathcal{G}_{\beta}$. Then:
  1. 1.
    $\tilde{\mathcal{F}}_{\lambda}$ is a normal filter on $\mathcal{G}_{\lambda}$ and contains every head pullback $\rho_{\beta,\lambda}^{-1}(H)$ for $\beta<\lambda$ and $H\in\mathcal{F}_{\beta}$.
  2. 2.
    $\tilde{\mathcal{F}}_{\lambda}$ is $\omega_{1}$–complete.
Proof.
By Definition 3.9, $\tilde{\mathcal{F}}_{\lambda}$ is (by construction) the smallest normal filter (or normal $\omega_{1}$–complete filter) containing the generator family $\mathcal{B}_{\lambda}$, so it is a normal filter and contains each $\rho_{\beta,\lambda}^{-1}(H)$.
For $\omega_{1}$–completeness: if $\mathrm{cf}(\lambda)=\omega$, closure under countable intersections is built into the defining criterion in Definition 3.9(2) by interleaving witnesses. If $\mathrm{cf}(\lambda)\geq\omega_{1}$, $\omega_{1}$–completeness follows from stage-bounding as proved in Remark 3.5. ∎
3.3 Hereditarily Symmetric Names in Countable Support Iterations
Having defined the limit filter $\tilde{\mathcal{F}}_{\lambda}$ (Definition 3.9) and verified its normality and $\omega_{1}$–completeness (Theorem 3.13), we now define the class of hereditarily symmetric names at stage $\lambda$ and verify the standard closure properties needed later for $\mathrm{ZF}$ and $\mathrm{DC}$ preservation.
Definition 3.14 (HS at stage $\lambda$).
Let $\langle\mathbb{P}_{\lambda},\mathcal{G}_{\lambda},\widetilde{\mathcal{F}}_{\lambda}\rangle$ be the symmetric system at stage $\lambda$. We write
$\mathrm{HS}_{\lambda}:=\mathrm{HS}(\mathbb{P}_{\lambda},\mathcal{G}_{\lambda},\widetilde{\mathcal{F}}_{\lambda})$
for the class of hereditarily $\widetilde{\mathcal{F}}_{\lambda}$-symmetric $\mathbb{P}_{\lambda}$-names, as in Definition 2.10. When $\lambda$ is fixed, we simply write $\mathrm{HS}$.
$\operatorname{sym}\!\bigl(\langle\dot{x}_{j}:j\in J\rangle\bigr)\;=\;\bigcap_{j\in J}\operatorname{sym}(\dot{x}_{j}).$ (1)
Justification of (1).
Let $\dot{s}=\langle\dot{x}_{j}:j\in J\rangle$ be the canonical sequence name. Then $g\in\operatorname{sym}(\dot{s})$ iff $g(\dot{s})=\dot{s}$, which (by the recursive definition of the action on names) holds iff $g(\dot{x}_{j})=\dot{x}_{j}$ for all $j\in J$, i.e. iff $g\in\bigcap_{j\in J}\operatorname{sym}(\dot{x}_{j})$. ∎
Theorem 3.15 (HS closure under standard name operations).
The class of HS names is closed under pairing, countable tuples, union, formation of subsets by bounded formulas (Separation instances over a fixed ambient set), and the usual forcing operations (e.g. check names).
Proof.
Pairing. If $\dot{x},\dot{y}$ are HS, then by the tuple–stabilizer identity (Equation (1))
$\operatorname{sym}\!\bigl(\langle\dot{x},\dot{y}\rangle\bigr)\;=\;\operatorname{sym}(\dot{x})\cap\operatorname{sym}(\dot{y}),$
and since $\tilde{\mathcal{F}}_{\lambda}$ is a filter (closed under finite intersections), the pair is HS.
Countable tuples. Let $\langle\dot{x}_{n}:n<\omega\rangle$ be a countable sequence of HS names, and let $\dot{s}$ denote the standard $\mathbb{P}_{\lambda}$–name for this tuple. By (1),
$\operatorname{sym}(\dot{s})\ =\ \bigcap_{n<\omega}\operatorname{sym}(\dot{x}_{n}).$
Since each $\operatorname{sym}(\dot{x}_{n})\in\tilde{\mathcal{F}}_{\lambda}$ and $\tilde{\mathcal{F}}_{\lambda}$ is $\omega_{1}$–complete (Theorem 3.13), we have $\operatorname{sym}(\dot{s})\in\tilde{\mathcal{F}}_{\lambda}$, so $\dot{s}$ is HS.
Union of an HS name. Let $\dot{A}$ be HS and define the usual union-name
$\dot{U}\ :=\ \Bigl\{\,\langle\tau,r\rangle:\exists\langle\sigma,p\rangle\in\dot{A}\ \exists\langle\tau,q\rangle\in\sigma\ (r\leq p,q)\Bigr\}.$
If $g\in\operatorname{sym}(\dot{A})$ then $g(\dot{U})=\dot{U}$ (because $g$ just transports witnesses in the defining formula), so $\operatorname{sym}(\dot{U})\supseteq\operatorname{sym}(\dot{A})\in\tilde{\mathcal{F}}_{\lambda}$. Hence $\dot{U}$ is HS.
Bounded Separation (over a fixed ambient set). Let $\dot{B}=\{z\in\dot{A}:\varphi(z,\bar{u})\}$ where $\dot{A}$ is HS and $\bar{u}$ is a tuple of HS parameters, and $\varphi$ is a bounded formula. If $g$ fixes $\dot{A}$ and $\bar{u}$, then by the Symmetry Lemma (Lemma 2.9), $g$ fixes $\dot{B}$. Hence
$\operatorname{sym}(\dot{B})\ \supseteq\ \operatorname{sym}(\dot{A})\cap\operatorname{sym}(\bar{u}),$
and the right-hand side lies in $\tilde{\mathcal{F}}_{\lambda}$ because $\tilde{\mathcal{F}}_{\lambda}$ is a filter (closed under finite intersections, and $\bar{u}$ is a finite tuple of parameters), so $\dot{B}$ is HS.
Check names and standard operations. If $a\in V$, then $\check{a}$ is fixed by all $g\in\mathcal{G}_{\lambda}$, so $\operatorname{sym}(\check{a})=\mathcal{G}_{\lambda}\in\tilde{\mathcal{F}}_{\lambda}$. Other usual name constructors are handled similarly; when a constructor uses only finitely many HS inputs, we use finite-intersection closure of $\tilde{\mathcal{F}}_{\lambda}$, and when it packages countably many HS inputs (e.g. countable tuples) we use (1) together with $\omega_{1}$–completeness (Theorem 3.13).
The hereditary clause in each case follows by induction on name rank. ∎
Remark 3.16 (Relation to finite-support symmetry).
In the finite-support setting (e.g. Karagila), filter closure under finite intersections suffices to obtain the HS closure properties. Our countable-support regime requires $\omega_{1}$-completeness to handle countable intersections of stabilizers arising from countably supported names; this is exactly where the Theorem 3.13 result is used.
3.4 $\mathrm{ZF}$ Preservation
Theorem 3.17 (ZF preservation for the symmetric submodel).
Assume we are working in the ambient background universe used for the iteration construction (cf. Remark 2.1). Let $V$ be the ground model for the forcing and assume $V\models\mathrm{ZF}$. Let $\lambda$ be a stage of the iteration and let $(\mathbb{P}_{\lambda},\mathcal{G}_{\lambda},\tilde{\mathcal{F}}_{\lambda})$ be the associated symmetric system, where $\tilde{\mathcal{F}}_{\lambda}$ is a normal $\omega_{1}$–complete filter of subgroups of $\mathcal{G}_{\lambda}$ (e.g. the limit filter from Theorem 3.13 in the countable–cofinality case). Let $G\subseteq\mathbb{P}_{\lambda}$ be VV–generic, and let $\mathrm{HS}$ denote the class of hereditarily $(\mathcal{G}_{\lambda},\tilde{\mathcal{F}}_{\lambda})$–symmetric $\mathbb{P}_{\lambda}$–names (Definition 3.14). Define
[…]
Then $M$ is a transitive inner model of $\mathrm{ZF}$ and
$V\ \subseteq\ M\ \subseteq\ V[G].$
Proof.
Write $\operatorname{sym}(\dot{x})=\operatorname{sym}_{\mathcal{G}_{\lambda}}(\dot{x})$ and let $\tilde{\mathcal{F}}_{\lambda}$ be the normal, $\omega_{1}$–complete limit filter from Theorem 3.13 (constructed at $\lambda$ as in Definition 3.9). We also use the tuple–stabilizer identity (1).
Transitivity and sandwich. If $a\in V$ then $\check{a}$ is fixed by all $g\in\mathcal{G}_{\lambda}$, hence HS and $a=\check{a}^{G}\in M$; thus $V\subseteq M$. Trivially $M\subseteq V[G]$. For transitivity, let $y\in M$ with $y=\dot{y}^{G}$ for some HS name $\dot{y}$. If $x\in y$ then $x=\sigma^{G}$ for some $(\sigma,p)\in\dot{y}$ with $p\in G$. Since $\dot{y}$ is hereditarily symmetric, every constituent $\sigma$ of $\dot{y}$ is HS, hence $x\in M$. Therefore $M$ is transitive.
Extensionality and Foundation. These are standard for symmetric submodels: $M$ is a transitive subclass of $V[G]$ with the inherited $\in$, and the forcing interpretation of names is extensional and well-founded; see, e.g., [2]. Thus $M$ satisfies Extensionality and Foundation.
Pairing and Union. If $\dot{x},\dot{y}$ are HS, then by (1) $\operatorname{sym}(\langle\dot{x},\dot{y}\rangle)=\operatorname{sym}(\dot{x})\cap\operatorname{sym}(\dot{y})\in\tilde{\mathcal{F}}_{\lambda}$, so the ordered pair is HS; hereditariness gives $\{\dot{x},\dot{y}\}$ HS as well.
For Union, let $\dot{A}$ be HS and define the usual union-name
$\dot{U}\ :=\ \Bigl\{\,\langle\tau,r\rangle:\exists\langle\sigma,p\rangle\in\dot{A}\ \exists\langle\tau,q\rangle\in\sigma\ (r\leq p,q)\Bigr\}.$
If $g\in\operatorname{sym}(\dot{A})$ then $g(\dot{U})=\dot{U}$ (since $g$ transports witnesses in the defining formula), so $\operatorname{sym}(\dot{U})\supseteq\operatorname{sym}(\dot{A})\in\tilde{\mathcal{F}}_{\lambda}$. Hence $\dot{U}$ is HS and therefore $U:=\dot{U}^{G}=\bigcup(\dot{A}^{G})$ belongs to $M$.
Infinity. $\check{\omega}$ is fixed by all automorphisms, hence HS; thus Infinity holds in $M$.
Separation. Let $A\in M$ be given by an HS name $\dot{A}$, fix HS parameters $\vec{\dot{p}}$, and let $\varphi$ be any formula. Define the name
$\dot{B}\ :=\ \{\ \langle\sigma,p\rangle\in\dot{A}\ :\ p\Vdash\varphi(\sigma,\vec{\dot{p}})\ \}.$
If $g\in\operatorname{sym}(\dot{A})\cap\operatorname{sym}(\vec{\dot{p}})$, then by the Symmetry Lemma (Lemma 2.9), $p\Vdash\varphi(\sigma,\vec{\dot{p}})\iff g(p)\Vdash\varphi(g\sigma,\vec{\dot{p}})$, whence $g\in\operatorname{sym}(\dot{B})$. Thus $\operatorname{sym}(\dot{B})\supseteq\operatorname{sym}(\dot{A})\cap\operatorname{sym}(\vec{\dot{p}})\in\tilde{\mathcal{F}}_{\lambda}$, so $\dot{B}$ is HS and $B:=\dot{B}^{G}=\{x\in A:\varphi^{M}(x,\vec{p})\}$. Hence $M$ satisfies Separation.
Power Set. Let $a\in M$ with $a=\dot{a}^{G}$ for HS $\dot{a}$. Fix $\alpha:=\operatorname{rank}(\dot{a})+1$ and let $\mathrm{Nm}_{\leq\alpha}$ be the (set-sized) collection in $V$ of $\mathbb{P}_{\lambda}$–names of rank $\leq\alpha$. Set
$\mathrm{HS}(\dot{a})\ :=\ \{\ \dot{x}\in\mathrm{Nm}_{\leq\alpha}\ :\ \Vdash\dot{x}\subseteq\dot{a}\ \text{ and }\ \dot{x}\in\mathrm{HS}\ \}.$
Then $\mathrm{HS}(\dot{a})$ is a set in $V$ (it is a definable subset of $\mathrm{Nm}_{\leq\alpha}$), and if $g\in\operatorname{sym}(\dot{a})$ then $g$ preserves both subsetting $\dot{a}$ and the HS property, so $\mathrm{HS}(\dot{a})$ is invariant under $\mathcal{G}_{\lambda}$. Consequently,
$\mathcal{P}^{M}(a)\ =\ \{\ \dot{x}^{G}:\dot{x}\in\mathrm{HS}(\dot{a})\ \}\ \in\ M,$
witnessed by the HS name coding $\mathrm{HS}(\dot{a})$ (using standard name-coding for sets of names). Therefore $M$ satisfies Power Set.
[…]
$\dot{\mathrm{rng}}\ :=\ \Bigl\{\ \bigl\langle\tau,p\bigr\rangle\ :\ \exists\langle\sigma,q\rangle\in\dot{A}\ \exists r\leq p,q\ \bigl\langle\langle\sigma,\tau\rangle,r\bigr\rangle\in\dot{f}\ \Bigr\}.$
If $g\in\operatorname{sym}(\dot{A})\cap\operatorname{sym}(\dot{f})$, then by definition $g\in\operatorname{sym}(\dot{\mathrm{rng}})$; hence $\operatorname{sym}(\dot{\mathrm{rng}})\supseteq\operatorname{sym}(\dot{A})\cap\operatorname{sym}(\dot{f})\in\tilde{\mathcal{F}}_{\lambda}$, so $\dot{\mathrm{rng}}$ is HS. By the forcing theorem and functionality of $\dot{f}$, we have
$\dot{\mathrm{rng}}^{G}\ =\ \{\,y\in V[G]:\exists x\in\dot{A}^{G}\ (\langle x,y\rangle\in\dot{f}^{G})\,\}\ =\ f``A.$
Thus $f``A\in M$, which is Collection and hence Replacement in $M$.
We have verified that $M$ is a transitive inner model of $\mathrm{ZF}$ with $V\subseteq M\subseteq V[G]$. ∎
Remark 3.18 (On the role of $\omega_{1}$-completeness).
In the countable-support setting, $\omega_{1}$-completeness of $\tilde{\mathcal{F}}_{\lambda}$ is needed to close HS under countable intersections of stabilizers (e.g., for closure under countable unions of HS constituents). If one proves Replacement by first enumerating the domain and setting $\dot{B}=\{\dot{y}_{n}:n<\omega\}$, then $\omega_{1}$-completeness is used to ensure $\bigcap_{n<\omega}\operatorname{sym}(\dot{y}_{n})\in\tilde{\mathcal{F}}_{\lambda}$.
In the proof of Theorem 3.17 given above, we instead use the relational range name, so the Replacement step itself only requires the finite intersection $\operatorname{sym}(\dot{A})\cap\operatorname{sym}(\dot{f})\in\tilde{\mathcal{F}}_{\lambda}$. Thus, for countable support the real novelty manifests earlier (in closure under countable unions and similar uses of $\omega_{1}$-completeness), rather than in the Replacement step per se.
3.5 Dependent Choice and Choice Fragments
Remark 3.19.
We use $\mathrm{DC}$ in the sense of Definition 2.11 (the set formulation).
Theorem 3.20 (DC Preservation (set-length stages)).
Assume $\lambda$ is a set ordinal and the ambient ground satisfies $\mathrm{ZFC}$. If $\tilde{\mathcal{F}}_{\lambda}$ is $\omega_{1}$-complete, then $M=V(\mathbb{P}_{\lambda})_{\mathrm{HS}}\models\mathrm{DC}$.
Proof.
Fix $G\subseteq\mathbb{P}_{\lambda}$ generic over $V$, and work in $V[G]$. Let $M=V(\mathbb{P}_{\lambda})_{\mathrm{HS}}$ be the symmetric model.
Let $A\in M$ and let $R\in M$ be a binary relation on $A$, and assume
$M\models\forall x\in A\ \exists y\in A\ R(x,y).$
Since $M$ is transitive (Theorem 3.17), the same statement holds in $V[G]$. Since $\mathbb{P}_{\lambda}$ is set forcing and $V\models\mathrm{ZFC}$, the forcing extension $V[G]\models\mathrm{ZFC}$, hence $V[G]\models\mathrm{DC}$. Therefore there exists a sequence $f:\omega\to A$ in $V[G]$ such that $R(f(n),f(n+1))$ for all $n$. Since $f$ is an $\omega$-sequence of elements of $M$, it suffices to show $M$ is closed under countable sequences.
For each $n<\omega$, we have $f(n)\in A\in M$, hence $f(n)\in M$ by transitivity. Choose HS names $\dot{a}_{n}\in\mathrm{HS}$ with $\dot{a}_{n}^{G}=f(n)$. By the countable-tuple closure of HS names (Theorem 3.15), the name $\langle\dot{a}_{n}:n<\omega\rangle$ is HS, and therefore $\langle f(n):n<\omega\rangle\in M$. This sequence witnesses $\mathrm{DC}$ in $M$. ∎
Remark 3.21 (Scope of DC obtained here).
The argument of Theorem 3.20 yields $\mathrm{DC}=\mathrm{DC}_{\omega}$ (the set formulation from Definition 2.11). Related preservation results for stronger fragments of dependent choice under $\kappa$-complete symmetry filters (under additional hypotheses on the forcing) appear, e.g., in [4, 1]. We do not claim any uncountable-length dependent choice (e.g. $\mathrm{DC}_{\omega_{1}}$) without additional hypotheses.
Corollary 3.22 (No infinite Dedekind-finite sets; no amorphous sets).
Under the hypotheses of Theorem 3.20, the symmetric model $M$ contains no infinite Dedekind-finite set and, in particular, no amorphous set.
Proof.
It is a theorem of $\mathrm{ZF}+\mathrm{DC}$ that every infinite set contains a countably infinite subset (see e.g. [2, Theorem 8.2]). Since $M\models\mathrm{DC}$ by Theorem 3.20, this applies inside $M$. An amorphous set—an infinite set with no infinite proper subset—would in particular be an infinite set with no countably infinite subset, contradicting $\mathrm{DC}$ in $M$. ∎
Remark 3.23 (Contrast with finite-support symmetric iterations).
In finite-support symmetric iterations, $\mathrm{DC}$ is not preserved in general at limit stages of cofinality $\omega$ (Remark 3.25), and amorphous or Dedekind-finite sets can appear in the resulting symmetric model. The classical example is Cohen’s first model: forcing with $\mathrm{Add}(\omega,\omega)$ and taking the symmetric submodel under finite-support permutations of coordinates yields a set of Cohen reals $A=\{a_{n}:n<\omega\}$ that is infinite but Dedekind-finite in the symmetric model, because no $\omega$-sequence enumerating $A$ is hereditarily symmetric (the enumeration stabilizer requires fixing all coordinates, falling outside any finite-support filter). See e.g. [2, Chapter 15] for details.
In contrast, the present countable-support framework ensures $\omega_{1}$-completeness of the limit filter at every cofinality-$\omega$ stage (Theorem 3.13), which forces HS closure under $\omega$-tuples (Theorem 3.15). The “forget the enumeration” mechanism of the finite-support setting is therefore unavailable: any countable enumeration of an infinite set is itself an HS name, and hence belongs to $M$.
Remark 3.24 (Choice landscape and filter closure).
Filter completeness governs which countable-sequence principles survive in the symmetric model:
  • If $\tilde{\mathcal{F}}_{\lambda}$ is $\omega_{1}$-complete (closed under countable intersections), then $\mathrm{HS}$ is closed under countable tuples, which yields $\mathrm{DC}$ in the symmetric model (Theorem 3.20).
  • Full $\mathrm{AC}$ is typically destroyed by the presence of enough symmetry.
In our countable-support framework the natural target is $\omega_{1}$-completeness, yielding DC while typically destroying full AC.
Remark 3.25 (Why finite support does not suffice here).
Countable support is used in two logically separate ways in this framework, distinguishing it from finite-support approaches (e.g., [3]).
(1) Countable sequences in the symmetric model. The proof of $\mathrm{DC}$ preservation (Theorem 3.20) does not rely on the forcing partial order being $\omega$-closed. Instead, the proof relies on $\omega_{1}$–completeness of the limit filter together with closure of $\mathrm{HS}$ under countable tuples (Theorem 3.15). Finite-support limit filters typically lack countable-intersection closure at limits of cofinality $\omega$, and therefore do not secure the tuple closure needed for $\mathrm{DC}$.
(2) Closure of symmetry under countable tuples. Our arguments repeatedly use that if $\langle\dot{x}_{n}:n<\omega\rangle$ consists of HS names, then the tuple itself is HS, which requires
$\bigcap_{n<\omega}\operatorname{sym}(\dot{x}_{n})\in\tilde{\mathcal{F}}_{\lambda}.$
This is exactly the $\omega_{1}$-completeness provided by Theorem 3.13.
(3) Applications requiring $\omega_{1}$-complete automorphism groups at limit stages. In the iterated unordered pairs model (Section 4), the impossibility of a choice function for $\mathcal{F}$ rests on each individual swap $\sigma_{\alpha}$ not belonging to the stabilizer of any choice-function name. For a DC-sequence through the pairs, by contrast, the stabilizer of the sequence name needs only to contain $\bigcap_{n<\omega}\operatorname{sym}(\dot{x}_{n})$, which $\omega_{1}$-completeness of the limit filter guarantees (Theorem 3.13). No individual $\sigma_{n}$ need fix the sequence name; $\omega_{1}$-completeness is what allows the tuple to be hereditarily symmetric as a whole. In a finite-support iteration the limit filter at a cofinality-$\omega$ stage is closed only under finite intersections of head pullbacks, which is insufficient: as shown explicitly in Proposition 4.4, a DC-counterexample exists in the finite-support model at stage $\omega$, and no HS name for the required sequence can be constructed. $\mathrm{DC}$ preservation is therefore an intrinsic consequence of $\omega_{1}$-completeness, not a byproduct of some other construction.
3.6 Iteration length
Proposition 3.26 (Set-length countable-support iterations).
Fix a countable-support iteration template as in Section 3. Then for every ordinal $\Theta$ the recursion defining
$\langle(\mathbb{P}_{\alpha},\mathcal{G}_{\alpha},\mathcal{F}_{\alpha}):\alpha\leq\Theta\rangle$
is well-defined. For each limit $\lambda\leq\Theta$, the induced limit filter $\tilde{\mathcal{F}}_{\lambda}$ is normal and $\omega_{1}$-complete, and the corresponding symmetric model $M_{\lambda}$ satisfies $\mathrm{ZF}$. If the ambient ground satisfies $\mathrm{ZFC}$, then $M_{\lambda}\models\mathrm{DC}$ as well.
Proof.
The successor step is the standard two-step symmetric-system construction (as cited in Definition 3.1). At limit stages $\lambda\leq\Theta$, we use Remark 3.2 to identify the countable-support limit with the direct limit when $\mathrm{cf}(\lambda)\geq\omega_{1}$, and we use the usual countable-support definition when $\mathrm{cf}(\lambda)=\omega$. In both cofinality cases, the limit filter is defined by Definition 3.9 and is normal and $\omega_{1}$-complete by Theorem 3.13. The symmetric model at stage $\lambda$ satisfies $\mathrm{ZF}$ by Theorem 3.17; and if the ambient ground is $\mathrm{ZFC}$, it satisfies $\mathrm{DC}$ by Theorem 3.20. ∎
Remark 3.27 (Class-length iterations).
We do not claim a general class-length $\mathrm{ZF}$-preservation theorem in this paper. The set-length results (Proposition 3.26, Theorems 3.17 and 3.20) are stated and proved for set-length iterations only, where the ambient forcing extension satisfies $\mathrm{ZFC}$ automatically.
At class length, $\mathrm{ZF}$ preservation for symmetric submodels requires hypotheses beyond pretameness: there are class forcings that are pretame yet whose symmetric submodels fail Collection or Power Set. Identifying the correct general hypothesis is a delicate open problem in class forcing theory.
Specific class-length symmetric iterations can often be analyzed directly by verifying each $\mathrm{ZF}$ axiom within the particular construction, as in [5]. The present framework supplies the limit-filter technology ($\omega_{1}$-completeness, Theorem 3.13) that such a verification would cite when checking closure of hereditarily symmetric names under countable tuples. We leave the formulation of a general class-length theorem to future work.
4 Application: The Iterated Unordered Pairs Model
We illustrate the framework with a self-contained construction that, for any uncountable cardinal $\kappa$ with $\mathrm{cf}(\kappa)\geq\omega_{1}$, produces a model of $\mathrm{ZF}+\mathrm{DC}+\neg\mathrm{AC}_{\kappa}(\mathcal{F})$, where $\mathcal{F}$ is a $\kappa$-indexed family of $2$-element sets of reals carrying no choice function. Each step of the iteration adds exactly one new unpickable pair, so the degree of $\mathrm{AC}$-failure is dialed by the iteration length. We also prove that the natural finite-support analogue fails to preserve $\mathrm{DC}$ at the first limit stage, showing that countable support is not merely sufficient but structurally necessary.
4.1 Setup: the step forcing and its symmetry
Fix an uncountable cardinal $\kappa$ with $\mathrm{cf}(\kappa)\geq\omega_{1}$ and a ground model $V\models\mathrm{ZFC}$. For each ordinal $\alpha<\kappa$ the step forcing is
$\dot{\mathbb{Q}}_{\alpha}\;=\;\mathrm{Add}(\omega)\times\mathrm{Add}(\omega),$
adding two independent Cohen reals, with the group $\dot{\mathcal{H}}_{\alpha}=\mathbb{Z}/2\mathbb{Z}$ acting by swapping coordinates: the nontrivial element $\sigma_{\alpha}$ sends $(s,t)\mapsto(t,s)$. Let $\dot{\mathcal{K}}_{\alpha}$ be the principal filter on $\dot{\mathcal{H}}_{\alpha}=\mathbb{Z}/2\mathbb{Z}$ generated by the whole group: the only filter element is $\mathbb{Z}/2\mathbb{Z}$ itself, i.e. $K\leq\mathbb{Z}/2\mathbb{Z}$ belongs to $\dot{\mathcal{K}}_{\alpha}$ if and only if $K=\mathbb{Z}/2\mathbb{Z}$.
After the generic at stage $\alpha$ is added, let $a_{\alpha},b_{\alpha}$ denote the two resulting Cohen reals. The individual reals $a_{\alpha}$ and $b_{\alpha}$ are not HS names: $\operatorname{sym}(\dot{a}_{\alpha})=\{\mathrm{id}\}\notin\dot{\mathcal{K}}_{\alpha}$, since $\{\mathrm{id}\}\neq\mathbb{Z}/2\mathbb{Z}$. The unordered pair
$\dot{P}_{\alpha}\;:=\;\{\dot{a}_{\alpha},\,\dot{b}_{\alpha}\}$
is an HS name: $\operatorname{sym}(\dot{P}_{\alpha})=\mathbb{Z}/2\mathbb{Z}\in\dot{\mathcal{K}}_{\alpha}$, because $\sigma_{\alpha}$ merely swaps the two elements of the pair and so fixes the pair as a set. Set
$\dot{\mathcal{F}}\;:=\;\bigl\langle\dot{P}_{\alpha}:\alpha<\kappa\bigr\rangle,$
the $\mathbb{P}_{\kappa}$-name for the function $\alpha\mapsto P_{\alpha}$, a $\kappa$-indexed family of $2$-element sets of reals.
Lemma 4.1 (Automorphisms used in the unordered-pairs argument).
In the countable-support iteration of $\kappa$ steps described above:
  1. 1.
    $\mathcal{G}_{\kappa}$ is abelian.
  2. 2.
    For each $\alpha<\kappa$, there exists an automorphism $g_{\alpha}\in\mathcal{G}_{\kappa}$ which acts by the nontrivial swap $\sigma_{\alpha}$ at stage $\alpha$ and acts trivially at every stage $\gamma\neq\alpha$.
  3. 3.
    For every $\beta\leq\alpha<\kappa$,
    $\rho_{\beta,\kappa}(g_{\alpha})=\mathrm{id}.$
    In particular, $g_{\alpha}\in\ker(\rho_{\beta,\kappa})$ whenever $\beta\leq\alpha$.
Proof.
For (1), argue by induction on $\delta\leq\kappa$ that $\mathcal{G}_{\delta}$ is abelian. At stage 0 this is trivial. At a successor stage, the new step group is $\mathbb{Z}/2\mathbb{Z}$, which is abelian, and the successor-stage action is coordinatewise, so automorphisms commute stage by stage. At a countable-cofinality limit, $\mathcal{G}_{\lambda}$ is an inverse limit of the earlier $\mathcal{G}_{\beta}$, hence abelian. At an uncountable-cofinality limit, $\mathcal{G}_{\lambda}$ is identified with $\bigcup_{\beta<\lambda}\mathcal{G}_{\beta}$ by Definition 3.1, hence again abelian.
For (2), fix $\alpha<\kappa$. At stage $\alpha+1$, the step group contributes the nontrivial swap $\sigma_{\alpha}$ exchanging the two Cohen coordinates added at stage $\alpha$. Extending by the identity at every later stage yields a coherent family of automorphisms, hence an element $g_{\alpha}\in\mathcal{G}_{\kappa}$ acting nontrivially only at stage $\alpha$.
For (3), if $\beta\leq\alpha$, then by construction $g_{\alpha}$ is trivial on every stage below $\beta$, so its restriction to stage $\beta$ is the identity. This is exactly the statement $\rho_{\beta,\kappa}(g_{\alpha})=\mathrm{id}$. ∎
4.2 Main theorem
Theorem 4.2 (Iterated Unordered Pairs Model).
Let $\kappa$ be an uncountable cardinal with $\mathrm{cf}(\kappa)\geq\omega_{1}$, and let $M_{\kappa}$ be the symmetric model obtained by the $\kappa$-step countable-support symmetric iteration described above. Then:
  1. 1.
    $M_{\kappa}\models\mathrm{ZF}+\mathrm{DC}$.
  2. 2.
    The indexed family $\mathcal{F}=\langle P_{\alpha}:\alpha<\kappa\rangle\in M_{\kappa}$ has no choice function in $M_{\kappa}$.
  3. 3.
    $M_{\kappa}\models\mathrm{ZF}+\mathrm{DC}+\neg\mathrm{AC}_{\kappa}(\mathcal{F})$.
Proof.
(1) $\mathrm{ZF}+\mathrm{DC}$. At each limit stage $\lambda\leq\kappa$, Theorem 3.13 gives a normal $\omega_{1}$-complete limit filter $\tilde{\mathcal{F}}_{\lambda}$. $\mathrm{ZF}$ preservation is Theorem 3.17. Since $\kappa$ is a set ordinal and $V\models\mathrm{ZFC}$, $\mathrm{DC}$ preservation is Theorem 3.20.
[…]
$\operatorname{sym}(\dot{\mathcal{F}})=\mathcal{G}_{\kappa}\in\tilde{\mathcal{F}}_{\kappa}.$
[…]
Now suppose for contradiction that $f\in M_{\kappa}$ is a choice function for $\mathcal{F}$, i.e. $f(\alpha)\in P_{\alpha}=\{a_{\alpha},b_{\alpha}\}$ for each $\alpha<\kappa$. Let $\dot{f}$ be an HS name for $f$, so $\operatorname{sym}(\dot{f})\in\tilde{\mathcal{F}}_{\kappa}$.
Since $\mathcal{G}_{\kappa}$ is abelian (Lemma 4.1(1)), conjugation in Definition 3.9(1) is vacuous. Hence membership $\operatorname{sym}(\dot{f})\in\tilde{\mathcal{F}}_{\kappa}$ is witnessed by finitely many head pullbacks: there exist $m<\omega$, stages $\beta_{0},\dots,\beta_{m-1}<\kappa$, and subgroups $H_{i}\in\mathcal{F}_{\beta_{i}}$ such that
$\bigcap_{i<m}\rho_{\beta_{i},\kappa}^{-1}(H_{i})\ \leq\ \operatorname{sym}(\dot{f}).$
Let $\beta^{*}=\max_{i<m}\beta_{i}<\kappa$. For any $\alpha\geq\beta^{*}$, Lemma 4.1(3) gives $g_{\alpha}\in\ker(\rho_{\beta_{i},\kappa})\leq\rho_{\beta_{i},\kappa}^{-1}(H_{i})$ for each $i<m$, hence $g_{\alpha}\in\operatorname{sym}(\dot{f})$.
Since $g_{\alpha}\in\operatorname{sym}(\dot{f})$, we have $g_{\alpha}(\dot{f})=\dot{f}$, so in the symmetric extension $g_{\alpha}(f(\alpha))=f(\alpha)$. But by Lemma 4.1(2), $g_{\alpha}$ applies the nontrivial swap at stage $\alpha$, exchanging $a_{\alpha}$ and $b_{\alpha}$: if $f(\alpha)=a_{\alpha}$, then $g_{\alpha}(a_{\alpha})=b_{\alpha}\neq a_{\alpha}$, and similarly if $f(\alpha)=b_{\alpha}$. This contradiction shows that no choice function for $\mathcal{F}$ exists in $M_{\kappa}$.
(3) Combines (1) and (2). ∎
Remark 4.3 (Each step does something; the length is the dial).
The symmetry argument in part (2) applies to each $\alpha$ independently: the $\mathbb{Z}/2\mathbb{Z}$ swap at stage $\alpha$ makes the pair $P_{\alpha}$ unpickable regardless of what happens at other stages. Running the iteration to any uncountable cardinal $\kappa^{\prime}\leq\kappa$ with $\mathrm{cf}(\kappa^{\prime})\geq\omega_{1}$ produces a model in which the truncated indexed family
$\langle P_{\alpha}:\alpha<\kappa^{\prime}\rangle$
has no choice function. By choosing $\kappa$ freely — any uncountable cardinal with $\mathrm{cf}(\kappa)\geq\omega_{1}$, with $\omega_{1}$ itself the simplest instance — one obtains $\mathrm{ZF}+\mathrm{DC}+\neg\mathrm{AC}_{\kappa}(\mathcal{F})$ at any such level. The restriction $\mathrm{cf}(\kappa)\geq\omega_{1}$ is used in the proof of part (2): the filter-membership witness for $\operatorname{sym}(\dot{f})$ then has finitely many head-pullback stages, all bounded below $\kappa$ by a single $\beta^{*}<\kappa$, allowing the tail-flip argument to go through. The case $\mathrm{cf}(\kappa)=\omega$ requires a separate argument tracking the cofinal sequence of witness stages and is not treated here.
4.3 Why finite support fails here
Proposition 4.4 (Finite support cannot preserve $\mathrm{DC}$ at the $\omega$-th limit).
Suppose one applies the same step forcing at each stage $\alpha<\omega$—that is, $\dot{\mathbb{Q}}_{\alpha}=\mathrm{Add}(\omega)\times\mathrm{Add}(\omega)$ with $\mathbb{Z}/2\mathbb{Z}$ symmetry—but uses a finite-support symmetric iteration. Then the resulting symmetric model $M^{\mathrm{fs}}_{\omega}$ fails $\mathrm{DC}$.
Proof.
Let $A:=\bigcup_{n<\omega}P_{n}$ be the union of all pairs in $M^{\mathrm{fs}}_{\omega}$ and define a binary relation on $A$ by
$x\mathrel{R}y\quad\Longleftrightarrow\quad\exists n<\omega\;(x\in P_{n}\;\wedge\;y\in P_{n+1}).$
The relation $R$ is total ($\forall x\in A\;\exists y\in A\;x\,R\,y$) and belongs to $M^{\mathrm{fs}}_{\omega}$ (its definition uses only the HS names $P_{n}$ and the successor function on $\omega$, both symmetric).
A $\mathrm{DC}$-sequence for $(A,R)$ would be an $\omega$-sequence $\langle x_{n}:n<\omega\rangle$ with $x_{n}\in P_{n}$ for each $n$, i.e., a partial choice function for $\{P_{n}:n<\omega\}$. Suppose $\dot{f}$ is an HS name for such a sequence in $M^{\mathrm{fs}}_{\omega}$. Then $\operatorname{sym}(\dot{f})$ lies in the finite-support limit filter $\tilde{\mathcal{F}}^{\mathrm{fs}}_{\omega}$ at stage $\omega$.
In the finite-support setting, every filter element at stage $\omega$ contains a subgroup that fixes all but finitely many coordinates: it is generated by finitely many head pullbacks $\rho_{k,\omega}^{-1}(H_{k})$ for $k<m$ and $H_{k}\in\mathcal{F}_{k}$. Hence for all but finitely many $n<\omega$, the swap automorphism $g_{n}$ (applying $\sigma_{n}$ at stage $n$, trivial elsewhere) belongs to $\operatorname{sym}(\dot{f})$.
But if $g_{n}\in\operatorname{sym}(\dot{f})$ then $g_{n}$ fixes $\dot{f}$, so $g_{n}(f(n))=f(n)$. Since $g_{n}$ swaps the two elements of $P_{n}$, this forces $f(n)$ to be fixed by the swap $a_{n}\leftrightarrow b_{n}$—impossible, as neither $a_{n}$ nor $b_{n}$ is individually HS. Contradiction; so no HS name for a DC-sequence through $(A,R)$ exists, and $M^{\mathrm{fs}}_{\omega}\not\models\mathrm{DC}$. ∎
Remark 4.5 (Why countable support resolves this).
In the countable-support iteration the limit filter $\tilde{\mathcal{F}}_{\omega}$ at stage $\omega$ is the smallest normal $\omega_{1}$-complete filter extending the head-generator family $\mathcal{B}_{\omega}$ (Definition 3.9(2) and Theorem 3.13). An HS name $\dot{f}$ for the DC-sequence only needs $\operatorname{sym}(\dot{f})\supseteq\bigcap_{n<\omega}\operatorname{sym}(\dot{x}_{n})$, and $\omega_{1}$-completeness guarantees this intersection lies in $\tilde{\mathcal{F}}_{\omega}$ (Theorem 3.15). No individual swap $\sigma_{n}$ need belong to $\operatorname{sym}(\dot{f})$; the tuple name is hereditarily symmetric as a whole, while each $P_{n}$ remains unpickable for the same reason as before. This is the precise sense in which countable support is not a convenience but a structural requirement.
5 Conclusion
We isolated the limit-stage filter construction needed for countable-support symmetric iterations built from the standard successor-step theory. The resulting limit filters are canonical, normal, and $\omega_{1}$-complete (Theorem 3.13). This $\omega_{1}$-completeness is the key input for closure of hereditarily symmetric names under countable tuples (Theorem 3.15), for the standard verification of $\mathrm{ZF}$ (Theorem 3.17), and, over a $\mathrm{ZFC}$ ground, for $\mathrm{DC}=\mathrm{DC}_{\omega}$ in the resulting symmetric model (Theorem 3.20).
We do not claim a general class-length $\mathrm{ZF}$-preservation theorem here. Specific class-length symmetric iterations may still be analyzable separately, but such results require additional hypotheses and care beyond the present set-length framework.
As a self-contained illustration, the iterated unordered pairs model (Section 4) demonstrates that for any uncountable cardinal $\kappa$ with $\mathrm{cf}(\kappa)\geq\omega_{1}$ the framework produces a model of $\mathrm{ZF}+\mathrm{DC}+\neg\mathrm{AC}_{\kappa}(\mathcal{F})$, with the degree of $\mathrm{AC}$-failure dialed by the iteration length. The same construction proves that finite support is structurally insufficient: the natural finite-support analogue fails $\mathrm{DC}$ at the first $\omega$-limit stage (Proposition 4.4).
In this sense the paper can be used modularly: once a concrete construction specifies its successor-stage symmetric systems, the limit-filter and $\mathrm{ZF}/\mathrm{DC}$ preservation arguments developed here may be cited separately, with any additional properties (e.g. chain conditions or cardinal behavior) verified in the specific application.
Acknowledgments. This work builds on the modern development of symmetric extensions and their iteration theory. We especially acknowledge Karagila’s foundational treatment of finite-support symmetric iterations [3], and the standard forcing background in Jech [2] and Kunen [7]. We also thank the authors of the other works cited herein. Any remaining errors are the author’s alone.
AI tools. Large language models (LLMs) were used as a writing aid for copy-editing, typesetting, and error-checking suggestions. All mathematical statements, proofs, and final editorial decisions were made by the author with whom final responsibility resides.